# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 4.2.4 Learning and cognitive capabilities
Psychometrics of learning and cognitive capabilities measures mental functions such as memory, reasoning, problem-solving, and comprehension to identify cognitive strengths and weaknesses. Recent studies adapt human psychometric tools to evaluate LLMs’ cognitive abilities. For example, Galatzer-Levy et al. [2024] use the Wechsler Adult Intelligence Scale (WAIS-IV) to show LLMs generally reach top human levels in verbal comprehension and working memory, though multimodal models lag in visual reasoning. Raven’s Progressive Matrices and other fluid intelligence tests reveal LLMs can match or surpass humans in analogical reasoning [Sartori and Orrú, 2023, Webb et al., 2022].
However, limitations persist. The Montreal Cognitive Assessment (MoCA) shows most LLMs display mild cognitive impairment, especially in visuospatial and executive tasks [Dayan et al., 2024]. Abstraction and Reasoning Challenge (ARC) further highlights deficits in abstract problem-solving [Wu et al., 2025]. Some works question the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], prompting new benchmarks and methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
Developmental perspectives also emerge: Wang et al. [2024e] leverage Piaget’s Theory of Cognitive Development to track human-like progression in LLMs, finding advanced models (e.g., GPT-4) reach cognitive abilities comparable to 20-year-old humans. Shah et al. [2024] identify developmental profiles in LLMs paralleling human development via psychometric theories like Cattell-Horn-Carroll.
@@
LLMs demonstrate strong performance on verbal comprehension, working memory, and analogical reasoning, often reaching or surpassing high human percentiles when evaluated with adapted psychometric tools like WAIS-IV and Raven’s Matrices [Galatzer-Levy et al., 2024, Webb et al., 2022]. However, they exhibit notable cognitive deficits, particularly on benchmarks like MoCA and ARC tasks [Wu et al., 2025, Dayan et al., 2024]. Several works introduce developmental and cognitive benchmarks (e.g., CogLM, CogBench) grounded in psychological theories to track reasoning and adaptability [Wang et al., 2024e, Coda-Forno et al., 2024]. While emergent reasoning appears at scale, challenges remain in interpretability, test validity, and generalization beyond surface-level statistical patterns [Li et al., 2024d, Löhn et al., 2024].
@@ 5 Psychometric evaluation methodology
This section examines the methodologies employed in LLM Psychometrics. As illustrated in Fig. 3, the methodological framework encompasses four key components: test formats (§ 5.1), data and task sources (§ 5.2), prompting strategies (§ 5.3), and model output and scoring (§ 5.4). The configuration of inference parameters (§ 5.5) is additionally discussed.
@@ 5.1 Test format
Personality traits Jiang et al. [2023], Serapio-García et al. [2023] Jiang et al. [2024b], Zheng et al. [2025] Frisch and Giulianelli [2024], Huang et al. [2024c]
Values Kovač et al. [2023], Zhong et al. [2024] Ren et al. [2024], Yao et al. [2025b] Chiu et al. [2025], Shen et al. [2024]
Morality Ji et al. [2024], Abdulhai et al. [2024] Sachdeva and van Nuenen [2025], Neuman et al. [2025] Chiu et al. [2025], Nunes et al. [2024]
Attitudes & opinions Bernardelle et al. [2024], Argyle et al. [2023] Röttger et al. [2024], Wright et al. [2024] -
Heuristics and biases Hagendorff et al. [2023], Yax et al. [2024] Healey et al. [2024], Wen et al. [2024a] Bai et al. [2024], Xie et al. [2024b]
Social interaction Sabour et al. [2024], Mittelstädt et al. [2024] Welivita and Pu [2024], Elyoseph et al. [2023] Zhou et al. [2024b], Wang et al. [2024d]
Psychology of language Duan et al. [2024a], Amouyal et al. [2024] - -
Learning and cognitive capabilities Wu et al. [2025], Coda-Forno et al. [2024] Galatzer-Levy et al. [2024] Lv et al. [2024]
Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations. Table 5 collects exemplary works illustrating the test formats used for each construct.
@@ 5.1.2 Unstructured tests
Unstructured testing often engages LLMs in open-ended, typically single-turn conversations that elicit specific constructs, mirroring real-world human-LLM interactions. For example, Jiang et al. [2024b] have LLMs generate long-form narratives to reveal personality traits. ValueBench uses advice-seeking queries to assess LLMs’ influence on users’ values [Ren et al., 2024], while Sachdeva and van Nuenen [2025] present moral dilemmas from Reddit, prompting LLMs to provide community-style moral judgments. To probe political attitudes, Röttger et al. [2024] convert PCT multiple-choice items into forced-choice and open-ended formats. For cognitive constructs, Healey et al. [2024] design pipelines to detect nuanced biases in free responses, and Welivita and Pu [2024] evaluate empathy by asking LLMs to respond to emotional scenarios. Many cognitive tests, such as WAIS-IV vocabulary and comprehension, also use open-ended questions [Galatzer-Levy et al., 2024].
@@
Advanced unstructured tests evaluate LLMs as agents in complex, contextualized role-playing scenarios, analyzing their decision-making in dynamic environments. Huang et al. [2024c] design LLM-Agents with distinct personalities and analyze them by replicating human-like behaviors in agentic simulations. Shen et al. [2024] measure values using four real-world vignettes: collaborative writing, education, public sectors, and healthcare. Chiu et al. [2025] present value and moral dilemmas to LLMs, examining their decisions and rationales. Cognitive construct tests are often more complex, extending to multi-agent settings; for example, Xie et al. [2024b], Bai et al. [2024] identify cognitive biases in multi-agent communication. SOTOPIA simulates open-ended social interactions among LLM agents to evaluate social intelligence [Zhou et al., 2024b, Wang et al., 2024d]. Lv et al. [2024] benchmark cognitive dynamics by having LLMs repeatedly complete cognitive questionnaires and reason over information flows.
@@ 5.2 Data and task sources
Personality traits Serapio-García et al. [2023], Barua et al. [2024] Ai et al. [2024], Jiang et al. [2023] Zeng [2024]
Values Fischer et al. [2023], Miotto et al. [2022] Shen et al. [2024], Meadows et al. [2024] Ye et al. [2025a], Moore et al. [2024]
Morality Abdulhai et al. [2024], Khandelwal et al. [2024] Hendrycks et al. [2021], Jin et al. [2022] Scherrer et al. [2023], Liu et al. [2024d]
Attitudes & opinions Argyle et al. [2023], Bernardelle et al. [2024] Durmus et al. [2023], Ceron et al. [2024] Wan and Chang [2024]
Heuristics and biases Hagendorff et al. [2023], Binz and Schulz [2023] Momennejad et al. [2023], Ando et al. [2023] Malberg et al. [2024], Xie et al. [2024b]
Social interaction [Ullman, 2023, Chen et al., 2024d] Shapira et al. [2023a], Street et al. [2024] Gandhi et al. [2023]
Psychology of language Cai et al. [2024], Duan et al. [2024a] He et al. [2024], Pérez-Mayos et al. [2021] -
Learning and cognitive capabilities Coda-Forno et al. [2024], Dayan et al. [2024] Song et al. [2024a], Wu et al. [2025] -
Data and task sources LLM Psychometrics can be 1) drawn from established psychometric inventories, 2) human-authored and custom-curated, and 3) synthesized by AI models. Table 6 provides an overview of the data and task sources for each construct.
@@
Established psychometric inventories offer standardized, well-validated tools for LLM evaluation. Common examples include BFI and HEXACO for personality [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes [Argyle et al., 2023, Bernardelle et al., 2024]; CRT for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; and False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d]. While these inventories enable straightforward, standardized assessment, their use with LLMs raises concerns about data contamination, response bias, and limited ecological validity for complex AI tasks [Ye et al., 2025a].
@@
Human-authored, custom-curated items offer tailored psychometric tests that are often more relevant to LLMs, enabling exploration of novel capability dimensions. For example, Shen et al. [2024] find SVS and PVQ insufficient for LLM value alignment, so they create 11 AI-informed value statements based on a systematic review of alignment literature. Ceron et al. [2024] highlight the unreliability of standard political questionnaires for LLMs and introduce tests using annotated voting-advice questionnaires from seven EU countries to assess stance consistency. For morality, Hendrycks et al. [2021] curate a dataset of over 130,000 open-world scenarios requiring moral judgments. In ToM, Soubki et al. [2024] develop a dataset from real spoken dialogues to address gaps between synthetic benchmarks and human behavior. While custom-curated items offer high ecological validity, their development and validation are labor-intensive, limiting scalability and diversity.
@@
Synthetic items—generated primarily by LLMs—are an emerging approach in LLM Psychometrics, enabling large-scale, diverse, and contextually rich test content. This method requires careful prompt engineering and validation. Some studies adapt established inventories to improve ecological validity, reduce data contamination, and enhance scalability. For instance, Ren et al. [2024] convert self-report items into advice-seeking queries to better reflect real-world human-AI interactions. Bhandari et al. [2025b] rewrite established items into semantically equivalent forms, validating them via semantic embedding similarity thresholds. Zhu et al. [2024a] diversify cognitive test items to enable more comprehensive analysis of LLM cognitive abilities. Other work generates synthetic tests entirely from scratch. Hadar-Shoval et al. [2025] demonstrate that LLM-generated cognitive tests correlate strongly with human performance on established tests. Ye et al. [2025a] create value-eliciting prompts and confirm the reliability and validity of the resulting measurements. Jiang et al. [2024a] introduce generative evolving testing, where an LLM-based Item Generator produces items of specified difficulty. Chiu et al. [2025] use GPT-4 to generate moral dilemmas for evaluating LLM moral judgment. For social intelligence, agentic simulations prompt LLMs to generate components of social interactions [Zhou et al., 2024b, Mou et al., 2024].
@@ 5.3 Prompting strategies
Role-playing prompts (persona or profiling prompts) embed demographic or personal attributes into the LLM context during testing. These prompts enable the generation of multiple simulated participants from a single LLM for statistical analysis. For example, Serapio-García et al. [2023] prepend persona instructions to each item to elicit diverse responses, allowing examining reliability and validity within one LLM. Similarly, Ye et al. [2025b] use value-anchoring prompts to create hundreds of LLM-based participants for value measurement, supporting the development of an LLM-specific value system. LLMs can represent a superposition of personalities, values, and perspectives [Kovač et al., 2023], prompting research into their adaptability. Role-playing prompts show that LLMs can shift between personality types, though steerability varies by model and dimension [Lu et al., 2023, Jiang et al., 2023, La Cava and Tagarelli, 2024, Li et al., 2024d]. Such prompts are also used to test the stability of LLM values [Kovač et al., 2024, Rozen et al., 2024] and to measure how well LLMs can mimic moral or political profiles and their susceptibility to related biases [Münker, 2024, Simmons, 2023, Wright et al., 2024]. Cognitive benchmarks, especially for social interactions, often use persona prompts in agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b], with task performance depending on the LLM’s ability to embody these roles. These prompts influence social-cognitive reasoning when controlling for other variables [Tan et al., 2024]. While role-playing capabilities enable diverse perspectives useful for social simulations, they can also lead to unstable values and perspectives, potentially compromising test validity and complicating value alignment.
@@
Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations, as recommended by Hagendorff [2023]. Chain of Thought (CoT) prompting [Wei et al., 2022] reduces biases [Hagendorff et al., 2023], enhances social intelligence [Shapira et al., 2023a], and improves cognitive capabilities [Coda-Forno et al., 2024]. Its variant Emotional CoT [Li et al., 2024e] boosts EI, while emotional prompts [Li et al., 2023a] enhance general cognitive abilities through emotional stimuli. Few-shot prompting [Brown et al., 2020] has demonstrated improvements in ToM performance [Moghaddam and Honey, 2023]. Research has also developed construct-specific prompting strategies. Zhou et al. [2024a] propose MFT-guided reasoning prompts for moral alignment, while Echterhoff et al. [2024], Sumita et al. [2024] focus on self-debiasing approaches. ToM-specific enhancements include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024]. Additionally, Zhao et al. [2025a] develop a framework using self-reflection prompts to examine both explicit and implicit social biases, where explicit bias measurement reflects implicit bias.
@@
Prompt perturbations can test the robustness of LLMs’ personality traits and cognitive abilities. Researchers have investigated whether LLMs can maintain stable personality, values, opinions, and cognitive abilities when item options are reordered [Schelb et al., 2025, Lee et al., 2024d], prompts are rephrased [Lee et al., 2024d, Fraser et al., 2022, Strachan et al., 2024a], prompt formats are altered [Moore et al., 2024, Schelb et al., 2025], or different languages are used [Moore et al., 2024, Cahyawijaya et al., 2024]. For instance, Faulborn et al. [2025] introduce 30 variations of prompts to evaluate the political biases of LLMs. In addition, Wen et al. [2024a] propose psychometric-inspired adversarial attacks to uncover implicit biases in LLMs. Similarly, Li et al. [2024d] utilize persuasive adversarial prompts (PAP) [Zeng et al., 2024a] to test the robustness of LLMs’ values. Since LLMs are sensitive to these perturbations, researchers have begun to scrutinize the reliability of the test results obtained under standard conditions [Ye et al., 2025a, Röttger et al., 2024, Dominguez-Olmedo et al., 2024, Strachan et al., 2024a].

### v3

@@ 4.2.4 Learning and cognitive capabilities
Psychometrics of learning and cognitive capabilities measures mental functions such as memory, reasoning, problem-solving, and comprehension to identify cognitive strengths and weaknesses. Early adaptations of human psychometric tools to LLMs reported striking successes. For instance, l1_galatzer2024cognitive used the Wechsler Adult Intelligence Scale (WAIS-IV) to show LLMs reaching top human levels in verbal comprehension, while other studies indicated that models could match or surpass humans in analogical reasoning on tests like Raven’s Progressive Matrices [l9_sartori2023language, l10_webb2022emergent].
However, a critical re-examination suggests these findings may be unreliable. High accuracy on such benchmarks often reflects “shortcut learning”, rather than genuine intelligence. Models exploit statistical patterns without engaging in true abstract reasoning. beger2025ai reveal that while models may achieve state-of-the-art results on the ARC-AGI benchmark, their solutions frequently rely on surface-level heuristics that fail to capture the intended abstractions. This brittleness is further exposed when models are tested for robustness and generalization: unlike humans, LLMs often fail to transfer analogical reasoning to novel domains or slightly modified tasks [lewis2024evaluatingrobustness, stevenson2024can].
From this perspective, documented failures on tests like the Montreal Cognitive Assessment (MoCA) [l3_dayan2024age] and the Abstraction and Reasoning Challenge (ARC) [l8_wu2025understanding, chollet2025arc] appear not as isolated deficits but as symptoms of a fundamental limitation in deep understanding. The “emergence” of reasoning capabilities is thus contested, with arguments that these traits may result from massive scale and memorization rather than novel cognitive properties [krakauer2025large]. Overall, LLMs are widely acknowledged to exhibit a “jagged intelligence,” characterized by uneven performance across tasks: while specific training and scaffolding have unlocked significant success in specialized domains such as mathematical Olympiads [luong2025advanced], these models often struggle with some tasks that could be simple and intuitive for humans.
@@
While early studies using adapted tools like WAIS-IV and Raven’s Matrices suggest LLMs possess human-like cognitive abilities [l1_galatzer2024cognitive, l10_webb2022emergent], a critical evaluation reveals these claims to be potentially misleading. Current methods often fail to account for shortcut learning, where high accuracy masks a lack of genuine abstract reasoning [beger2025ai]. The cognitive abilities of LLMs are shown to be brittle, collapsing under task modification or transfer, which starkly contrasts with human cognition [lewis2024evaluatingrobustness, stevenson2024can]. Deficits on benchmarks like MoCA and ARC highlight a fundamental lack of robust, generalizable understanding. Overall, models demonstrate a “jagged intelligence,” achieving high performance in specialized domains (e.g., mathematics and coding) through targeted training and scaffolding [luong2025advanced, novikov2025alphaevolve], while failing on tasks that are trivial and intuitive for humans. This reveals a fundamentally uneven cognitive profile compared to human intelligence. Future psychometric frameworks should move beyond accuracy-based metrics to assess the robustness, process, and validity of reasoning [mitchell2025artificial].
@@ 5 Psychometric evaluation methodology
This section examines the methodologies employed in LLM Psychometrics. As illustrated in Fig.˜3, the methodological framework encompasses four key components: test formats (§˜5.1), data and task sources (§˜5.2), prompting strategies (§˜5.3), and model output and scoring (§˜5.4). The configuration of inference parameters (§˜5.5) is additionally discussed.
@@ 5.1 Test format
Personality traits jiang2023evaluating_and_inducing, serapio2023personality_traits_in_llms jiang2024personallm, zheng2025lmlpa frisch2024llm_agents_in_interaction, huang2024designing
Values kovac2023llms_as_superpositions, zhong2024cultural ren2024valuebench, yao2025clave chiu2025dailydilemmas, shen2024valuecompass
Morality ji2024moralbench, abdulhai2024moral sachdeva2025normative, neuman2025analyzing chiu2025dailydilemmas, nunes2024large
Attitudes & opinions bernardelle2024mapping, argyle2023_out_of_one rottger2024political, wright2024revealing -
Heuristics and biases hagendorff2023reasoning_bias, yax2024studying healey2024evaluating, wen2024evaluating bai2024fairmonitor, xie2024mindscope
Social interaction sabour2024emobench, si1_mittelstadt2024large welivita2024large, elyoseph2023chatgpt si3_zhou2023sotopia, si4_wang2024sotopia
Psychology of language pl1_duan2024hlb, pl24_amouyal2024large - -
Learning and cognitive capabilities l8_wu2025understanding, coda2024cogbench l1_galatzer2024cognitive lv2024coggpt
Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations. Table˜4 collects exemplary works illustrating the test formats used for each construct.
@@ 5.1.2 Unstructured tests
Unstructured testing often engages LLMs in open-ended, typically single-turn conversations that elicit specific constructs, mirroring real-world human-LLM interactions. For example, jiang2024personallm have LLMs generate long-form narratives to reveal personality traits. ValueBench uses advice-seeking queries to assess LLMs’ influence on users’ values [ren2024valuebench], while sachdeva2025normative present moral dilemmas from Reddit, prompting LLMs to provide community-style moral judgments. To probe political attitudes, rottger2024political convert PCT multiple-choice items into forced-choice and open-ended formats. For cognitive constructs, healey2024evaluating design pipelines to detect nuanced biases in free responses, and welivita2024large evaluate empathy by asking LLMs to respond to emotional scenarios. Many cognitive tests, such as WAIS-IV vocabulary and comprehension, also use open-ended questions [l1_galatzer2024cognitive].
@@
Advanced unstructured tests evaluate LLMs as agents in complex, contextualized role-playing scenarios, analyzing their decision-making in dynamic environments. huang2024designing design LLM-Agents with distinct personalities and analyze them by replicating human-like behaviors in agentic simulations. shen2024valuecompass measure values using four real-world vignettes: collaborative writing, education, public sectors, and healthcare. chiu2025dailydilemmas present value and moral dilemmas to LLMs, examining their decisions and rationales. Cognitive construct tests are often more complex, extending to multi-agent settings; for example, xie2024mindscope, bai2024fairmonitor identify cognitive biases in multi-agent communication. SOTOPIA simulates open-ended social interactions among LLM agents to evaluate social intelligence [si3_zhou2023sotopia, si4_wang2024sotopia]. lv2024coggpt benchmark cognitive dynamics by having LLMs repeatedly complete cognitive questionnaires and reason over information flows.
@@ 5.2 Data and task sources
Personality traits serapio2023personality_traits_in_llms, barua2024on_the_psychology_of_gpt4 ai2024is_self_knowledge, jiang2023evaluating_and_inducing zeng2024quantifying
Values fischer2023does, miotto2022who_is_gpt3 shen2024valuecompass, meadows2024localvaluebench ye2025gpv, moore2024large
Morality abdulhai2024moral, khandelwal2024moral hendrycks2021aligning, jin2022make scherrer2023evaluating, liu2024evaluating
Attitudes & opinions argyle2023_out_of_one, bernardelle2024mapping durmus2023towards, ceron2024beyond wan2024white
Heuristics and biases hagendorff2023reasoning_bias, binz2023using momennejad2023evaluating, ando2023evaluating malberg2024comprehensive_cognitive_biases, xie2024mindscope
Social interaction [tom1_ullman2023large, tom14_chen2023tobench] tom4_shapira2023clever, tom9_street2024llms tom5_gandhi2023understanding
Psychology of language cai2024large, pl1_duan2024hlb pl9_he2024large, pl11_perez2021assessing -
Learning and cognitive capabilities coda2024cogbench, l3_dayan2024age l11_song2024m3gia, l8_wu2025understanding -
Data and task sources LLM Psychometrics can be 1) drawn from established psychometric inventories, 2) human-authored and custom-curated, and 3) synthesized by AI models. Table˜5 provides an overview of the data and task sources for each construct.
@@
Established psychometric inventories offer standardized, well-validated tools for LLM evaluation. Common examples include BFI and HEXACO for personality [serapio2023personality_traits_in_llms, barua2024on_the_psychology_of_gpt4]; SVS, PVQ, WVS, and VSM for values [fischer2023does, miotto2022who_is_gpt3]; MFT and DIT for morality [abdulhai2024moral, khandelwal2024moral]; ANES and PCT for political attitudes [argyle2023_out_of_one, bernardelle2024mapping]; CRT for heuristics and biases [hagendorff2023reasoning_bias, binz2023using]; and False-Belief Tasks for ToM [tom1_ullman2023large, tom14_chen2023tobench]. While these inventories enable straightforward, standardized assessment, their use with LLMs raises concerns about data contamination, response bias, and limited ecological validity for complex AI tasks [ye2025gpv].
@@
Human-authored, custom-curated items offer tailored psychometric tests that are often more relevant to LLMs, enabling exploration of novel capability dimensions. For example, shen2024valuecompass find SVS and PVQ insufficient for LLM value alignment, so they create 11 AI-informed value statements based on a systematic review of alignment literature. ceron2024beyond highlight the unreliability of standard political questionnaires for LLMs and introduce tests using annotated voting-advice questionnaires from seven EU countries to assess stance consistency. For morality, hendrycks2021aligning curate a dataset of over 130,000 open-world scenarios requiring moral judgments. In ToM, tom19_soubki2024views develop a dataset from real spoken dialogues to address gaps between synthetic benchmarks and human behavior. While custom-curated items offer high ecological validity, their development and validation are labor-intensive, limiting scalability and diversity.
@@
Synthetic items—generated primarily by LLMs—are an emerging approach in LLM Psychometrics, enabling large-scale, diverse, and contextually rich test content. This method requires careful prompt engineering and validation. Some studies adapt established inventories to improve ecological validity, reduce data contamination, and enhance scalability. For instance, ren2024valuebench convert self-report items into advice-seeking queries to better reflect real-world human-AI interactions. bhandari2025evaluating rewrite established items into semantically equivalent forms, validating them via semantic embedding similarity thresholds. zhu2024dynamic diversify cognitive test items to enable more comprehensive analysis of LLM cognitive abilities. Other work generates synthetic tests entirely from scratch. hadar2025feasibility demonstrate that LLM-generated cognitive tests correlate strongly with human performance on established tests. ye2025gpv create value-eliciting prompts and confirm the reliability and validity of the resulting measurements. jiang2024raising introduce generative evolving testing, where an LLM-based Item Generator produces items of specified difficulty. chiu2025dailydilemmas use GPT-4 to generate moral dilemmas for evaluating LLM moral judgment. For social intelligence, agentic simulations prompt LLMs to generate components of social interactions [si3_zhou2023sotopia, si8_mou2024agentsense].
@@ 5.3 Prompting strategies
Role-playing prompts (persona or profiling prompts) embed demographic or personal attributes into the LLM context during testing. These prompts enable the generation of multiple simulated participants from a single LLM for statistical analysis. For example, serapio2023personality_traits_in_llms prepend persona instructions to each item to elicit diverse responses, allowing examining reliability and validity within one LLM. Similarly, ye2025generative_psycho_lexical use value-anchoring prompts to create hundreds of LLM-based participants for value measurement, supporting the development of an LLM-specific value system. LLMs can represent a superposition of personalities, values, and perspectives [kovac2023llms_as_superpositions], prompting research into their adaptability. Role-playing prompts show that LLMs can shift between personality types, though steerability varies by model and dimension [lu2023illuminating, jiang2023evaluating_and_inducing, lacava2024open_models_closed_minds]. Such prompts are also used to test the stability of LLM values [kovavc2024stick, rozen2024llms] and to measure how well LLMs can mimic moral or political profiles and their susceptibility to related biases [munker2024towards, simmons2023moral, wright2024revealing]. Cognitive benchmarks, especially for social interactions, often use persona prompts in agentic simulations [si3_zhou2023sotopia, si8_mou2024agentsense, huang2023emotionbench], with task performance depending on the LLM’s ability to embody these roles. These prompts influence social-cognitive reasoning when controlling for other variables [tan2024phantom]. While role-playing capabilities enable diverse perspectives useful for social simulations, they can also lead to unstable values and perspectives, potentially compromising test validity and complicating value alignment.
@@
Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations, as recommended by hagendorff2023machine. Chain of Thought (CoT) prompting [wei2022chain] reduces biases [hagendorff2023reasoning_bias], enhances social intelligence [tom4_shapira2023clever], and improves cognitive capabilities [coda2024cogbench]. Its variant Emotional CoT [li2024enhancing] boosts EI, while emotional prompts [li2023emotion_prompt] enhance general cognitive abilities through emotional stimuli. Few-shot prompting [brown2020language] has demonstrated improvements in ToM performance [tom8_rahimi2023boosting]. Research has also developed construct-specific prompting strategies. zhou2024rethinking propose MFT-guided reasoning prompts for moral alignment, while echterhoff2024cognitive, sumita2024cognitive focus on self-debiasing approaches. ToM-specific enhancements include SymbolicToM [tom10_sclar2023minding] and temporal decomposition-based prompting [tom26_sarangi2025decompose, tom16_hou2024timetom]. Additionally, zhao2025explicit develop a framework using self-reflection prompts to examine both explicit and implicit social biases, where explicit bias measurement reflects implicit bias.
@@
Prompt perturbations can test the robustness of LLMs’ personality traits and cognitive abilities. Researchers have investigated whether LLMs can maintain stable personality, values, opinions, and cognitive abilities when item options are reordered [schelb2025ru, lee2024do_llms_have_distinct], prompts are rephrased [lee2024do_llms_have_distinct, fraser2022does, tom6_strachan2024testing], prompt formats are altered [moore2024large, schelb2025ru], or different languages are used [moore2024large, cahyawijaya2024high]. For instance, faulborn2025only introduce 30 variations of prompts to evaluate the political biases of LLMs. In addition, wen2024evaluating propose psychometric-inspired adversarial attacks to uncover implicit biases in LLMs. Similarly, li2024quantifying_ai_psychology utilize persuasive adversarial prompts (PAP) [zeng2024johnny] to test the robustness of LLMs’ values. Since LLMs are sensitive to these perturbations, researchers have begun to scrutinize the reliability of the test results obtained under standard conditions [ye2025gpv, rottger2024political, dominguez2024questioning, tom6_strachan2024testing].